OpenStudy (anonymous):

Arithmetic Sequence A man puts $300 into a savings account at the end of each year for 15 yrs. If his money draws 3% simple interest, how much will he have in the account at the end of 15 yrs?

OpenStudy (math&ing001):

Lets break it down : At the end of 1st year: 3% interest out of $0 = $0 total = 0+300 = $300 At the end of 2nd year: 3% interest out of $300 = $9 total = 9+300+300 = $609 At the end of 3rd year: 3% interest out of $609 = $18.27 total = 18.27+609+300 = $927.27 So the general term would be: u(n)=3%*u(n-1)+u(n-1)+300=(0.03+1)*u(n-1)+300 u(n) = 1.03*u(n-1) + 300

OpenStudy (anonymous):

@math&ing001 is the answer 426? just making sure

OpenStudy (math&ing001):

I'm sorry, but it couldn't be $426 because at the end of the 3rd year he already has $927.27 and it's only going to increase. It's gonna be much more at the 15th year.

OpenStudy (math&ing001):

First thing we could do is make u(n) only dependent of 'n' so we can find u(15) directly. You can use v(n) = 1.03^n * 300 as geometric sequence And find its sum, meaning u(15) = v(1) + v(2) + ... + v(15) Did you study the sum of a geometric sequence ?

OpenStudy (anonymous):

hmm no. But the problem stated that I should use the arithmetic sequence